# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

### v1

@@ III Methodology
All-in-One image restoration (AiOIR) methods have gained significant attention for their ability to address multiple types of degradations within a unified model. These methods are particularly valuable in real-world scenarios where images may be simultaneously affected by various artifacts, such as noise, blur, and adverse weather conditions like fog and rain. In this section, we provide a comprehensive analysis of AiOIR methods from multiple perspectives. Firstly, we undertake a review of both novel and traditional network architectures and deliberate on the related approaches for their implementation (classifying these methods into several typical network designs). Subsequently, we further investigate diverse learning strategies, encompassing training methodologies and perspectives, to augment the accuracy of image reconstruction and enhance the overall restoration quality. Finally, there exist other techniques that further optimize AiOIR models, including Prompt Learning, Mixture-of-Experts (MoE), Multimodal Model, and so on.
@@ III-A Network Design for AiOIR
Existing all-in-one image restoration (AiOIR) methods, as a form of Multi-Task Learning (MTL), employ various architectural designs to handle inputs and outputs across multiple tasks, enabling efficient information sharing among them. Despite significant differences in their architectures, AiOIR methods can be categorized into ten representative frameworks, as illustrated in Fig. 3 and described below.
(a) Task-Specific Encoders and Decoders: This straightforward approach assigns each type of degradation to a specifically designed encoder-decoder pair. For example, one encoder-decoder might handle low-light enhancement, while another addresses image denoising. This setup requires prior knowledge of the degradation type to select the appropriate components. However, in real-world scenarios, images often suffer from multiple or unknown degradations, making this approach less practical.
@@
(b) Shared Decoder with Multiple Heads: Models in this category share a common decoder but have multiple heads trained individually for different degradation types. For instance, the All-in-One model by Li et al. [56] processes various types of adverse weather images using shared weights, requiring separate training for each degradation type.
(c) Unified Encoder-Decoder Architecture: These models use a single encoder-decoder architecture without separate heads or tails, designed to remove one specific type of noise at a time. They serve as the foundation for prompt-based methods like TransWeather [64], AirNet [63], and TANet [109].
(d) Shared Backbone with Multiple Decoders: Universal models adopt mixed inputs without any task-specific indicators, using a shared backbone for feature extraction and multiple task-specific decoders. For example, BIDeN [101] follows this approach. However, this reintroduces the complexity of multiple decoders and requires extensive supervision from degradation labels.
@@
(f) Mixture-of-Experts (MoE) Architecture: In MoE-based models, inputs are routed to different experts via a gating mechanism. MEASNet [110], for instance, considers both pixel-level and global features—including low-frequency and high-frequency components—to select the appropriate expert for image restoration.
(g) Pretrained Model Prior: Models like Perceive-IR [111] and DA-CLIP [76] utilize frozen vision-language models, such as CLIP [112], DINO [113], and DINO-v2 [114]. These models predict high-quality feature embeddings to enhance the restoration process by leveraging semantic alignment between vision and language.
(h) Visual Prompting: These models employ a single encoder-decoder architecture and inject learnable visual prompts at multiple decoding stages to implicitly predict degradation conditions. Examples include works by Potlapalli et al.[65], Li et al.[72], and Fan et al. [115]. The prompts guide the decoder in adaptively recovering various degraded images, serving as lightweight, plug-and-play modules with minimal additional parameters.
(i) Textual and Multimodal Prompting: Extending the concept of prompting, models like Yan et al.[94], Conde et al.[62], and Tian et al. [116] incorporate textual or multimodal prompts. These models allow for natural language instructions or combine visual and textual cues to steer the restoration process, enhancing adaptability to unknown degradations.
@@
AiOIR models have evolved from task-specific architectures requiring prior knowledge of degradation types to more flexible and unified frameworks capable of handling multiple degradations without explicit information. A significant trend in recent research is the incorporation of prompt-based techniques—visual, textual, or multimodal—that guide the restoration process and enhance adaptability. By leveraging prompts, models can effectively handle unknown degradations and offer user-controllable restoration, making prompt technology a promising direction in AiOIR.
Figure 4: Transfer learning scenarios in AiOIR.
@@ III-B Learning Strategies of AiOIR
Beside network design (as discussed in Sec. III-A), robust learning strategies are crucial for achieving satisfactory results in AiOIR. In this section, we discuss several promising learning strategies in the field. We begin with continual learning, exploring how it can prevent catastrophic forgetting. Next, we focus on contrastive learning and its application in complex degradation scenarios, emphasizing its role in enhancing the discriminative power of image features. Then, we highlight the potential of Multi-Task Learning (MTL) to optimize performance across various degradation tasks, pointing out the importance of addressing task relationships and conflicts. Finally, we introduce the concept of machine unlearning, exploring its potential for privacy protection.
@@ III-B1 Continual Learning
AiOIR methods can be categorized into two learning styles from the perspective of transfer learning: MTL and sequential learning. MTL involves learning different tasks simultaneously, while sequential learning involves learning different tasks in sequence. An illustration of transfer learning scenarios [117, 118] is shown in Fig. 4. In human cognition, forgetting is a gradual process; except in rare cases, people do not suddenly lose their memory. However, in computational models, catastrophic forgetting [119] often occurs, where after learning new knowledge, the model almost completely forgets previously learned content. In the field of AiOIR, since a single network is expected to restore images with multiple degradations, the model must learn knowledge related to various degradations, making catastrophic forgetting more likely. To enable the model to incrementally accumulate knowledge and avoid catastrophic forgetting, researchers have proposed novel learning strategies, such as review learning [120] and sequential learning [121]. These methods are inspired by continual learning [122, 123] and replace mixed training (mixing datasets with different degradations) with sequential training. It is worth noting that the learning order of multiple tasks is crucial to the quality of image restoration. In MiOIR [121], the authors studied the effect of training order on the results and pointed out the impact of multi-task orders. In SimpleIR [120], the authors delve into the entropy difference distribution for various IR tasks. They propose determining the training dataset order based on abnormally high loss values and the intrinsic difficulty of tasks, measured by the entropy difference between original and degraded images.
Figure 5: Schematic illustration of Contrastive Learning in AiOIR.
@@ III-B2 Contrastive Learning
One of the significant challenges in image restoration is effectively handling unseen tasks and degradation types. The vast variability in possible degradations can severely impede a model’s generalization capabilities, making it less effective when confronted with new, unseen data. To address this issue, researchers have drawn inspiration from contrastive learning techniques that have proven successful in both high-level and low-level tasks [124, 125, 126]. Contrastive learning approaches typically serve as an additional form of regularization to improve the generalization of single-task restoration models [127, 128, 126, 129, 130]. By incorporating contrastive regularization, these methods aim to enhance model performance across a wide array of image restoration applications. In contrastive learning, the definitions of positive and negative samples can be flexibly adjusted (as illustrated in Fig.5), allowing researchers to tailor the learning process to better suit specific tasks and datasets. This flexibility ultimately enhances the model’s adaptability and performance in diverse image restoration scenarios[111]. Moreover, contrastive learning-based loss functions have been proposed to obtain discriminative degradation representations in the latent space [63, 66, 72], further improving the model’s ability to distinguish between different types of degradations and to generalize to unseen ones.
III-B3 Multi-Task Learning
@@
MTL is a learning paradigm that leverages shared representations and knowledge across tasks, allowing models to learn more efficiently and effectively. By jointly learning from multiple objectives, MTL can improve generalization, reduce overfitting, and achieve better performance on individual tasks. It is widely used in various domains [131, 132, 133], where related tasks can benefit from shared information. However, in the context of AiOIR, the optimization process often receives less attention, leading to an oversight of the complex relationships and potential conflicts between multiple degradations in mixed training scenarios. Unlike unified models (e.g.[65, 121, 64, 63]) that employ mixed training for multiple-degradation restoration, some novel studies approach AiOIR from the perspective of MTL to address the inconsistencies and conflicts among different degradations, treating each degradation as an independent task. By focusing on the optimization process and the interactions between tasks, these methods aim to mitigate conflicts and enhance overall performance. As illustrated in Fig. 6, we can broadly classify AiOIR methods in MTL into two types: task grouping and task balancing.
Figure 6: Schematic comparison of multi-task learning methods in AiOIR with traditional all-in-one methods.
@@
Task Balancing. Conversely, the work by Art [135] proposes a straightforward yet effective loss function that incorporates task-specific reweighting. This method dynamically balances the contributions of individual tasks, fostering harmony among diverse tasks. By adjusting the weight of each task based on its specific characteristics and performance, this approach aims to mitigate conflicts and enhance the overall effectiveness of the restoration process.
III-B4 Machine Unlearning
Privacy protection is an increasingly critical issue in the field of artificial intelligence (AI), particularly as models become more integrated into sensitive applications. One promising solution to address this challenge is machine unlearning [136]. This innovative approach aims to effectively erase the influence of private data from trained models, allowing them to function as if that sensitive information was never utilized during the training process.
Existing methodologies for unlearning can be broadly categorized into two types: exact unlearning [137], which seeks to completely eliminate the impact of specific data points, and approximate unlearning [138], which aims to diminish the influence of such data to a certain degree. While the concept of unlearning has been extensively explored in various contexts, including classification tasks and federated learning scenarios, its application in the realm of end-to-end image restoration remains largely uncharted.
To fill this gap, Su et al. [139] propose an innovative framework that applies machine unlearning techniques to AiOIR models. In their work, they define a scenario in which certain types of degradation data—such as haze or rain—are treated as private information that must be effectively ”forgotten” by the model. Importantly, this process is designed to preserve the model’s performance on other types of degradations, ensuring that the overall functionality remains intact. To achieve this goal, the authors introduce a technique called Instance-wise Unlearning, which leverages adversarial examples in conjunction with gradient ascent methods. This approach not only enhances the model’s ability to forget specific data but also maintains its robustness across various image restoration tasks.
@@ III-C Key Improvements of AiOIR
In addition to the network design in Sec. III-A and learning strategies in Sec. III-B, there are other key techniques aiming at improving AiOIR models. In this section, we preliminarily divided the key improvements of AiOIR Model into the following three areas, i.e., Prompt Learning, Mixture-of-Experts (MoE), and Multimodal Model, which correspond to Sec. III-C1, Sec. III-C2, Sec. III-C3, respectively. We also illustrate some other improvements, including deep unfolding methods, masked image modeling, etc. For the sake of clarity, we have listed the representative works on all-in-one models based on the three types in Fig. 7.
Figure 7: The overview of AiOIR methods. This figure categorizes AiOIR into three types based on their key improvements, namely the Prompt-based models (indicated with a blue background), MoE-based models (indicated with an orange background), Multimodal models (indicated with a green background).
@@ III-C1 Prompt Learning
Prompt learning, initially successful in Natural Language Processing (NLP) [140, 141, 142], aims to tap into the knowledge that the language model itself possesses by providing instructions or relevant information. Inspired by the ability of effectively modeling task-specific context, prompts have been used for finetuning to vision tasks [143, 144]. To be specific, unlike single-task image restoration, learnable prompts enable better parameter-efficient adaptation of models facing with multiple degradations. Recently, various prompts have been explored in AiOIR, as an adaptive lightweight module to encode degradation representation in the restoration network. The core idea is to enable pre-trained models to better understand and perform downstream tasks by constructing visual, textual or multimodal prompts, which are as follows.
Visual Prompting. Visual prompting has been extensively explored in various studies [145, 146, 80, 61], addressing both high-level and low-level vision problems, which has sparked significant interest across different visual domains. In the context of AiOIR, prompt learning-based techniques have been implemented, allowing a generic model to automatically select prompts for specific tasks, resulting in excellent performance. Among these, vision-based prompts are the most widely used. For instance, AirNet [63] utilizes degradation encoder representations to guide network recovery, while Transweather [64] applies queries to direct the recovery process. An increasing number of researchers are applying the concept of prompt learning to develop AiOIR models.
PromptIR [65] is one of the most representative works, integrating a prompt block into the U-Net architecture [147] to enhance AiOIR. This prompt block takes prompt components and the output from the previous transformer block as inputs, with its output fed into the next transformer block. The prompts serve as an adaptive, lightweight plug-and-play module as shown in Fig. 8, encoding degradation context across multiple scales in the restoration network. In contrast, ProRes [61] introduces a target visual prompt added to an input image, resulting in a “prompted image”. This prompted image is then flattened into patches, with the weights of ProRes frozen, while learnable prompts are randomly initialized for new tasks or datasets. Building on these ideas, PromptGIP [80] proposes a training method similar to masked autoencoding [148], where certain portions of question and answer images are randomly masked. This prompts the model to reconstruct these patches from the unmasked areas, and during inference, input-output pairs are assembled as task prompts to facilitate image restoration. PIP [72] introduces a novel Prompt-In-Prompt learning framework for universal image restoration, which employs two innovative prompts: a degradation-aware prompt and a basic restoration prompt. PIP relies solely on the input image to recover a clean image, without requiring any prior knowledge of the degradation present in the image.
@@
Textual Prompting. As elaborated above, many studies leverage learnable prompts to guide restoration. Nevertheless, these models primarily concentrated on learning visual prompts based on distributional differences in the training data. Due to the presence of a semantic gap [149], accurately identifying degradation types remained a challenge, resulting in only moderate restoration performance. In this context, several innovative studies [116, 62, 70, 94, 74, 150] have introduced textual prompts, advancing the field of AiOIR. For example, TextPromptIR [94] utilizes a task-specific fine-tuned BERT [151] to accurately understand textual instructions and generating semantic prompts for concerned all-in-one tasks. Unlike traditional methods, text prompts also allow users to describe specific degradations in their images using natural language, which the model then interprets to apply appropriate restoration techniques. This approach enhances the adaptability and usability of AiOIR, making them more accessible to non-expert users and enabling the models to generalize better across different types of degradation scenarios. As GPT-3.5 [152] becomes increasingly popular, Clarity ChatGPT [150], combined with advanced visual models, offers users a simple and efficient way to perform complex image manipulation and enhancement through natural language interaction. Furthermore, AutoDIR [70] can handle images with unknown degradations in unseen task via an intuitive text prompt, empowering users to refine restoration results in alignment with their visual preferences.
Multimodal Prompting. Recently, vision-language models (VLMs) have demonstrated the great potential of applying pretrained VLMs to improve downstream tasks with generic visual and text representations. A traditional VLM typically includes a text encoder and an image encoder, aiming to learn aligned multimodal features from noisy image-text pairs using contrastive learning. This greatly facilitates multimodal prompts in textual for holistic representation and visual for multiscale detail representation. MPerceiver [71], a groundbreaking multimodal prompt learning method, is designed to harness the generative priors of Stable Diffusion [153], enhancing the adaptiveness, generalizability, and fidelity of AiOIR. Both prompts are dynamically refined based on predictions from the CLIP image encoder, allowing for adaptive handling of various unknown degradations. More recently, DA-CLIP [112] leveraged pre-trained large-scale vision models to excel in AiOIR with an image controller that detects degradation and adapts the fixed CLIP image encoder to generate high-quality content embeddings from corrupted inputs.

### v2

@@ III Methodology
All-in-one image restoration methods have gained significant attention for their ability to address multiple types of degradation within a unified model. These methods are particularly valuable in real-world scenarios where images may be simultaneously affected by various artifacts, such as noise, blur, and adverse weather conditions. This section presents a comprehensive analysis of AiOIR methods from multiple perspectives. Firstly, we examine both novel and traditional network architectures, discussing their implementation approaches and classifying them into several typical network designs. Subsequently, we further investigate diverse learning strategies, which encompass training methodologies and perspectives, to improve image reconstruction accuracy and overall restoration quality. Finally, we examine other advanced techniques that optimize AiOIR models, including Prompt Learning, Mixture-of-Experts (MoE), Multimodal Model, and Agentic IR System. In Fig. 2, we present the representative works on AiOIR according to the timeline and method types.
@@ III-A Network Design for AiOIR
Existing all-in-one image restoration methods, as a form of multi-task learning (MTL), employ various architectural designs to handle inputs and outputs across multiple tasks, enabling efficient information sharing among them. Despite significant differences in their architectures, AiOIR methods can be broadly categorized into ten representative frameworks, as illustrated in Fig. 3 and described below.
(a) Task-Specific Encoders and Decoders: This straightforward approach assigns a specifically designed encoder-decoder pair to each type of degradation. For example, one encoder-decoder might handle low-light enhancement, while another addresses image denoising. This setup requires prior knowledge of the degradation type to select the appropriate components. However, in real-world scenarios, images often suffer from multiple or unknown degradations, making this approach less practical.
@@
(b) Shared Decoder with Multiple Heads: Models in this category share a common decoder but have multiple heads trained individually for different degradation types. For instance, Li et al. [53] proposed a method for processing various types of adverse weather images using shared weights, but requiring separate training for each degradation type.
(c) Unified Encoder-Decoder Architecture: These models use a single encoder-decoder architecture without separate heads or tails, often serving as foundational designs that can be adapted for multi-degradation removal. There are methods like TransWeather [29], AirNet [54], TANet [78], and AdaIR [79].
(d) Shared Backbone with Multiple Decoders: Universal models adopt mixed inputs without any task-specific indicators, using a shared backbone for feature extraction and multiple task-specific decoders. For example, BIDeN [80] follows this paradigm. However, this reintroduces the complexity of multiple decoders and requires extensive supervision with degradation labels.
@@
(f) Mixture-of-Experts Architecture: In MoE-based models, inputs are routed to different experts via a gating mechanism. For instance, MEASNet [82] considers both pixel-level and global features (including low-frequency and high-frequency components) to select the appropriate expert for image restoration.
Figure 3: Ten categories of the image restoration method prototype (the last nine are AiOIR methods) (a) independent encoders and decoders. (b) multiple heads and a shared decoder. (c) a single branch end-to-end model. (d) a shared encoder and multiple decoders. (e) a reusable pretrained middle-level backbone. (f) the mixture-of-experts architecture. (g) pretrained large vision-language models to guide image restoration. (h) visual prompts to guide image restoration. (i) textual prompts or instructions to guide image restoration. (j) a question answering paradigm.
(g) Pretrained Model Prior: Models like DINO-IR [83], Perceive-IR [84], and DA-CLIP [85] utilize frozen vision-language models (e.g., CLIP [86], DINO [87], and DINO-v2 [88]). These models predict high-quality feature embeddings to enhance the restoration process by leveraging semantic alignment between vision and language.
(h) Visual Prompting: These models employ a single encoder-decoder architecture and inject learnable visual prompts at multiple decoding stages to implicitly predict degradation conditions. Examples include works like [58, 89, 90, 91, 92]. The prompts guide the decoder in adaptively recovering various degraded images, serving as lightweight plug-and-play modules with minimal additional parameters.
(i) Textual and Multimodal Prompting: Extending the concept of prompting, models like [93, 94, 95] incorporate textual or multimodal prompts. These models allow for natural language instructions or combine visual and textual cues to steer the restoration process, enhancing adaptability to unknown degradations.
@@
AiOIR models have evolved from task-specific architectures requiring prior knowledge of degradation types to more flexible and unified frameworks capable of handling multiple degradations without explicit information. Moreover, recent methods increasingly adopt hybrid designs, which we further categorize by their component labels. For example, (g+h) methods such as [97] combine DepthAnything [98] with learnable visual prompts; (f+h) methods like [82] use task‐specific prompts to guide multi-expert selection; (f+i) methods like [99] leverage the power of pretrained vision-language models to enrich the diversity of weather-specific knowledge based on an MoE architecture; (h+i) methods, with representative works including [84, 100], embed both visual and textual prompts into a unified framework. With this concise labeling, emerging hybrids can be directly slotted into our taxonomy.
@@ III-B Learning Strategies of AiOIR
Besides network design (as discussed in Sec. III-A), robust learning strategies are crucial for achieving satisfactory results in AiOIR. In this section, we discuss several promising learning strategies in this field. We begin with continual learning, exploring how it can prevent catastrophic forgetting. Next, we focus on contrastive learning and its application in complex degradation scenarios, emphasizing its role in enhancing the discriminative power of image features. We then discuss the zero-shot learning in AiOIR, emphasizing its ability to generalize to unseen degradations. Additionally, we highlight the potential of multi-task learning (MTL) to optimize performance across various degradation tasks, pointing out the importance of addressing task relationships and conflicts. Finally, we introduce the concept of machine unlearning, exploring its potential for privacy protection.
@@ III-B1 Continual Learning
AiOIR methods can be categorized into two learning styles from the perspective of transfer learning [101, 102]: multi-task learning and sequential learning. MTL involves learning different tasks simultaneously, while sequential learning involves learning different tasks in sequence. In learning models, catastrophic forgetting [103] often occurs, where after learning new knowledge, the model almost completely forgets previously learned content. In the field of AiOIR, since a single network is expected to restore images with multiple degradations, the model must learn knowledge related to various degradations, making catastrophic forgetting more likely. To enable the model to incrementally accumulate knowledge and avoid catastrophic forgetting, researchers proposed novel learning strategies, such as review learning [104] and sequential learning [105]. These methods are inspired by continual learning [106, 107] and replace mixed training (mixing datasets with different degradations) with sequential training. It is worth noting that the learning order of multiple tasks is crucial for the quality of IR. In MiOIR [105], the authors study the effect of training order on the results and point out the impact of multi-task orders. In SimpleIR [104], the authors delve into the entropy difference distribution for various IR tasks. They propose determining the training dataset order based on abnormally high loss values and the intrinsic difficulty of tasks, measured by the entropy difference between original and degraded images. Recently, UniCoRN [108] topologically sorts both single and synthetic multi-degradation datasets based on a parent-child relationship graph and, further sort datasets within a level in the descending order of size to minimize the effect of catastrophic forgetting.
@@ III-B2 Contrastive Learning
One of the significant challenges in IR is effectively handling unseen tasks and degradation types. The vast variability in possible degradations can severely impede a model’s generalization capabilities, making it less effective when confronted with new, unseen data. To address this issue, researchers have drawn inspiration from contrastive learning techniques that have proven successful in both high-level and low-level tasks [109, 110, 111]. Contrastive learning approaches typically serve as an additional form of regularization to improve the generalization of single-task restoration models [112, 113, 111, 114, 115]. By incorporating contrastive regularization, these methods aim to enhance model performance across a wide array of image restoration applications. In contrastive learning, the definitions of positive and negative samples can be flexibly adjusted, allowing researchers to tailor the learning process to better suit specific tasks and datasets. This flexibility ultimately enhances the model’s adaptability and performance in diverse image restoration scenarios [84]. Moreover, contrastive learning-based loss functions have been proposed to obtain discriminative degradation representations in the latent space [54, 57, 89, 116], further improving the AiOIR model’s ability to distinguish between different types of degradations and to generalize to unseen ones.
III-B3 Zero-shot Learning
By restoring images with distortions that are absent in the training data, zero-shot image restoration tackles the challenge of generalizing effectively to unseen degradation types. For instance, methods (e.g.,  [117, 118, 119, 120]) leverage pretrained diffusion models as generative priors, seamlessly integrating degraded images as conditions into the sampling process. In the case of AiOIR, MPerceiver [121] demonstrates strong zero-shot and few-shot capabilities across six unseen tasks. Additionally, TAO [122] employs test-time adaptation for AiOIR tasks, achieving results comparable to or better than traditional supervised methods. The core difficulty in these scenarios lies in the distribution shift between training and test data. This requires models to generalize beyond their training experience and adapt to previously unseen degradation types. Test-time adaptation (TTA) (e.g.,  [123, 124, 125]) has emerged as a key technique in addressing this issue, allowing models to dynamically adjust their parameters during the testing phase to better align with the characteristics of the degraded input images. By enabling the model to restore images with unknown degradations, zero-shot AiOIR aims to provide a more robust and versatile solution, making it a crucial step towards real-world applicability.
III-B4 Multi-task Learning
@@
MTL is a learning paradigm that leverages shared representations and knowledge across tasks, allowing models to learn more efficiently and effectively. By jointly learning from multiple objectives, MTL can improve generalization, reduce overfitting, and achieve better performance on individual tasks. Traditional MTL typically involves a combination of semantic tasks (e.g., classification, detection, segmentation). The tasks exhibit significant semantic differences and also possess distinct output spaces. Typical MTL architectures share a backbone and then branch into task-specific heads, while optimization techniques such as uncertainty weighting, GradNorm, or gradient surgery address conflicts between task gradients [126, 127, 128]. AiOIR, by contrast, treats each image degradation as a separate task but operates on the same input–output domain (i.e., pixel-level image reconstruction). Losses remain largely reconstruction-focused (perceptual loss, pixel loss), and task conflicts arise at the gradient level, often manifesting as mutual interference among degradation models. The design of the degradation-aware module is often insufficient to resolve conflicts among different degradations. The optimization process often receives less attention, leading to an oversight of the complex relationships and potential conflicts between multiple degradations in mixed training scenarios. Unlike unified models (e.g.,  [58, 105, 29, 54]) that employ mixed training, some novel studies address AiOIR from the perspective of MTL to resolve the inconsistencies and conflicts among different degradations, by treating each degradation as an independent task. By focusing on the optimization process and the interactions between tasks, these methods aim to mitigate conflicts and enhance overall performance. We can broadly classify AiOIR methods in MTL into two types: task grouping and task balancing.
@@
Task Balancing. Conversely, Wu et al. [130] proposed a straightforward yet effective loss function that incorporates task-specific reweighting. This method dynamically balances the contributions of individual tasks, fostering harmony among diverse tasks. By adjusting the weight of each task based on specific characteristics and performance, this approach mitigates conflicts and enhances the overall effectiveness. Subsequently, Wu et al. proposed TUR [131], a task-aware optimization strategy that introduces adaptive task-specific regularization for AiOIR.
III-B5 Machine Unlearning
Privacy protection is an increasingly critical issue in the field of artificial intelligence, particularly as models become more integrated into sensitive applications. One promising solution to address this challenge is machine unlearning [132]. This innovative approach aims to effectively erase the influence of private data from trained models, allowing them to function as if that sensitive information was never utilized during the training process. Existing methodologies for unlearning can be broadly categorized into two types: exact unlearning [133], which seeks to completely eliminate the impact of specific data points, and approximate unlearning [134], which aims to diminish the influence of such data to a certain degree. While the concept of unlearning has been extensively explored in various contexts, including classification tasks and federated learning scenarios, its application in the realm of end-to-end IR remains largely uncharted.
To fill this gap, Su et al. [135] proposed an innovative framework that applied machine unlearning techniques to AiOIR models. They defined a scenario in which certain types of degradation are treated as private information that must be effectively forgotten by the model. Importantly, this process is designed to preserve the model’s performance on other types of degradations, ensuring that the overall functionality remains intact. To achieve this goal, the authors introduced a technique called instance-wise unlearning, which leveraged adversarial examples in conjunction with gradient ascent methods. This approach not only enhances the model’s ability to forget specific data but also maintains its robustness across various image restoration tasks.
@@ III-C Key Improvements of AiOIR
In addition to the network design in Sec. III-A and learning strategies in Sec. III-B, there are other key techniques to improve AiOIR models. In this section, we preliminarily categorize the key improvements of AiOIR models into the following four areas, i.e., Prompt Learning, Mixture-of-Experts (MoE), Multimodal Model and Agentic IR System, which correspond to Sec. III-C1, Sec. III-C2, Sec. III-C3, Sec. III-C4, respectively. We also illustrate some other improvements, including deep unfolding methods, mask image modeling, hypernetworks, etc. For the sake of clarity, we list the representative works on all-in-one models based on the four types in Fig. 4.
Figure 4: The overview of AiOIR methods. This figure categorizes AiOIR into four types based on their key improvements, namely the Prompt-based models, MoE-based models, Multimodal models and Agentic IR systems.
@@ III-C1 Prompt Learning
Prompt learning, initially successful in natural language processing (NLP) [136, 137, 138], aims to tap into the knowledge that is inherent in the language model by providing instructions or relevant information. Inspired by the ability to effectively model task-specific context, prompts have been used for fine-tuning for vision tasks [139, 140]. To be specific, unlike single-task image restoration, learnable prompts enable better parameter-efficient adaptation of models facing multiple degradations. Recently, various prompts have been explored in AiOIR, as an adaptive lightweight module to encode degradation representation in the network. The core idea is to enable pretrained models to better understand and perform downstream tasks by constructing visual, textual or multimodal prompts, which are as follows.
Visual Prompting. Visual prompting has gained widespread attention for tackling both high-level and low-level vision tasks [141, 142, 143]. In AiOIR, prompt learning enables models to adaptively select task-specific prompts, achieving strong performance across diverse degradations. For example, AirNet [54] leverages degradation-aware features, while Transweather [29] uses query-based guidance. This growing trend highlights the increasing integration of prompt learning in AiOIR research.
PromptIR [58] is one of the most representative works, integrating a prompt block into the U-Net architecture to enhance Restormer [50]. The prompts serve as an adaptive, lightweight plug-and-play module as shown in Fig. 5, enabling multi-scale degradation-aware guidance across layers. Extending this idea, ProRes [59] embeds visual prompts directly into the input image. PromptGIP [96] proposes a training method similar to masked autoencoding [144], where certain portions of question and answer images are randomly masked. This prompts the model to reconstruct these patches from the unmasked areas. PIP [89] further introduces a novel Prompt-In-Prompt learning framework for AiOIR, which employs two innovative prompts: a degradation-aware prompt and a basic restoration prompt. To improve task-specific representation, Wen et al. [143] propose multi-axis prompt and multi-dimension fusion network for all-in-one traffic weather removal by incorporating prompts from three axes. CPL [116] establishes a new paradigm that extends contrastive learning principles from feature-level discrimination to the previously unexplored domain of prompt-level alignment. Recently, Wu et al. [145] introduce prompt distribution learning, which models prompts as distributions rather than fixed parameters.
@@
Textual Prompting. While earlier AiOIR studies focused on learnable visual prompts, their effectiveness was limited by the semantic gap [146], making it difficult to accurately identify degradation types. To address this, recent works [95, 94, 61, 93, 99, 147] explore textual prompts for more precise and user-friendly guidance. For instance, TextPromptIR [93] leverages a fine-tuned BERT model to interpret task-specific instructions, enabling semantic understanding of degradations. Text prompts allow users to describe degradations in natural language, enhancing model adaptability and accessibility. Building on this, Clarity ChatGPT [147] and AutoDIR [61] combine large language models with visual features to support intuitive and interactive IR, even in unseen tasks. Extending to hyperspectral image restoration, PromptHSI [74] introduces text-guided, frequency-aware modulation, showcasing the growing potential of language-driven AiOIR.
Multimodal Prompting. Vision-language models (VLMs) have shown strong potential in AiOIR by leveraging aligned visual and textual representations. Contrastive learning allows VLMs to train image and text encoders, thereby enabling prompts that integrate text for holistic semantics and vision for fine-grained details. Building on this, MPerceiver [121] utilizes stable diffusion priors and CLIP-based dynamic prompts to adaptively handle diverse degradations. DA-CLIP [86] introduces an image controller to align degraded inputs with high-quality CLIP embeddings. CyclicPrompt [148] fuses CLIP-derived weather knowledge, conditional vectors, and textual prompts for context-aware restoration. MP-HSIR [75] focuses on HSI restoration by integrating spectral, visual, and textual prompts. Recently, VLU-Net [149] leverages CLIP-derived high-dimensional features for automatic selection of degradation-aware keys in deep unfolding networks. These works highlight multimodal prompting in enhancing adaptability and generalization in AiOIR.